OpenStudy (anonymous):

how would you solve the equation (x+a)(x+b)=23 when (a) and (b) are real numbers and what are the solutions

OpenStudy (amistre64):

x^2 + (a+b)x + ab -23

OpenStudy (amistre64):

-(a+b)/2 +- sqrt((a+b)^2 - 4(ab - 23))/2

OpenStudy (amistre64):

do you need actual numbers for the solution?

OpenStudy (amistre64):

integer of otherwise?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

x+b = 23 or x+a = 23....would be the only integer solutions I see

OpenStudy (amistre64):

if we are not bound to integers, then there cold be alot of solutions

OpenStudy (amistre64):

x^2 +(a+b)x +____ = (23 -ab) x^2 + (a+b)x + (a+b)^2/4 = 23 - ab + (a+b)^2/4 (x+(a+b)/2)^2 = 23 -ab +(a+b)^2/4

OpenStudy (amistre64):

all we have to go on really is that they y intercept is (ab-23) and I dont see a 'one' solution to this

OpenStudy (amistre64):

any thoughts?

OpenStudy (anonymous):

nothing but thank you